OpenStudy (anonymous):

f(x) = 3sin(2x), my amplitude is 3, is my period 2?

Nnesha (nnesha):

\[\huge\rm y = A \sin (B \theta-C )+ D\] here a is amplitude and period should be 2pi/B

jimthompson5910 (jim_thompson5910):

the amplitude should be |A| to make sure it's never negative

Nnesha (nnesha):

yeah amplitude never going to be negative ;)

OpenStudy (xapproachesinfinity):

the period is found using this \[2x=2\pi \Longrightarrow x=\pi\]

OpenStudy (anonymous):

so the period is pi

OpenStudy (xapproachesinfinity):

yes

Nnesha (nnesha):

APPLE pi

OpenStudy (xapproachesinfinity):

hahaha :)

OpenStudy (xapproachesinfinity):

the period in general is obtained by: \[period=\frac{2\pi}{|B|}\]

OpenStudy (anonymous):

i drew a graph of the function and put the period as 2 pi and it was marked right, now I am totally confused :(